OpenStudy (deplorableboi):

Please help solve 1 variable equation!!! 3(x + 1) = −2(x − 1) + 6

OpenStudy (sapphiremoon):

I assume we're solving for x. First, distribute the coefficients. \[3x + 3 = -2x + 2 + 6\] Add any constants that are on the same side. \[3x + 3 = -2x + 8\] Add 2x to both sides to move that term to the LHS. \[3x + 2x + 3 = -2x + 2x + 8\] \[5x + 3 = 8\] Subtract 3 from both sides to get both constants on the RHS. \[5x + 3 - 3 = 8 - 3\] \[5x = 5\] Divide both sides by 5 (the coefficient of x). \[x = 1\]

OpenStudy (mhchen):

You need to isolate the variable by itself. Any ideas on how to do that?